I have a 55 gallon tank. I would like a 15% sulfuric acid, 85% water solution in my tank. I have 98% sulfuric acid in quarts. How much sulfuric acid and how much water will I need to add to my take to have a 15% sulfuric acid solution?
Answer by richwmiller(17219)

from water point of view
.85*55=.02a+b,
a+b=55
a=8.41837 gallons of acid , b=46.5816 gallons of water
from acid point of view
.15*55=.98x
x = 8.41837 gallons of acid
55-x=46.5816 gallons of water